# Radiation-dominated polar emitting region of an accreting X-ray pulsar – I. Polarization- and spectrum-dependent structure, and the emergent continuum

### v1

@@ 3.2 Results of the numerical calculations
The boundaries of the colour bars related to the 2D distributions indicate the minimum and maximum values approached within the shown domains (for the velocity, these values fall within the corresponding boundary intervals). The difference between $\theta_{0}$, $\theta_{1}$ and $\theta_{2}$ in calculations and their values indicated above lies within the grid steps. Seediness at the front of the shock wave in the electron temperature distributions is the trace of the colour interpolation and is neither a manifestation of the scheme instabilities nor a property of the grid functions.
@@
It follows from the description above that, in simulations, the lower accretion rate, the lower absolute values of PF are reached (Figs 13). That is, unsaturated Comptonization of the radiation of the ordinary mode can result in an emergent spectrum comparably close to the spectrum of the extraordinary mode (at least at $\dot{M}_{17}\lesssim 1$, Figs 1, 4, 5 and 6). It is especially clear in the case of the near-critical value of the accretion rate $\dot{M}_{17}=0.2$ ($r_{*}(r_{\rm up})=1.15$, $L_{\rm X,\,37}\approx 0.7$), see Fig. 6. The calculated PF in the range 112020 keV is in a satisfactory agreement with theoretical that of Burnard et al. (1991) and with characteristic observational values (e.g. Tsygankov et al. 2022; Long et al. 2023).
@@ 3.4 Effects of bulk Comptonization
In the absence of a magnetic field, first-order bulk Comptonization would be the reason of a shift in the maximum of the emergent spectrum toward higher energies relative to the characteristic energy $2.8kT_{0}$. This is a typical picture (see Psaltis & Lamb 1999, fig. 1). In a strong magnetic field this effect takes place as well. The emergent spectrum of the extraordinary mode is shifted to higher energies relative to the value of $kT_{0}$ mainly due to the influence of first-order bulk Comptonization, while the thermal effect of the shock wave is not very significant for the shift. The spectrum of the ordinary mode can be affected by the heating of the shock wave. If a high-energy hump in the emergent spectrum of the ordinary mode is formed, it usually corresponds to the energy $\sim 2kT_{\rm s}$, where $T_{\rm s}$ is the maximum electron temperature (approaching in the shock wave).
@@
The influence of the $m_{\rm e}v^{2}/3$ term can be illustrated by varying the velocity at the upper boundary. Consider a filled funnel. In Fig. 5 the structure, the emergent spectra and the polarization fraction are shown for $v(r_{\rm up})=1.3\times 10^{10}$ cm s-1 ($\dot{M}_{17}=1$, $L_{\rm X,\,37}\approx 1.7$). The decrease in $v(r_{\rm up})$ leads to the shift of the high-energy cutoff toward low energies, as expected. A considerable decrease in $v(r_{\rm up})$ (up to $\sim c/3$) demands a decrease in $\gamma$ for the existence of solutions in the framework of the model. Although such solutions contain a significantly cooler shock (up to $T_{\rm s}\approx 9$ keV at $T_{0}\approx 7$ keV for $\dot{M}_{17}=3$), the relatively softer spectral tail remains nevertheless rather hard, and the PF distribution is similar to that given in Fig. 2. The possibility of the distinction of $v(r_{\rm up})$ from the free-fall velocity was mentioned by Poutanen et al. (2013).
@@ 4.1 The model
The overall picture of matter breaking in the radiation-dominated regime is thus as follows. Five zones can be distinguished in the accretion channel above the magnetic pole: (1) the cold accretion flow falls near-freely in a radiation field of relatively low energy densities; (2) the flow begins to fluently brake in the shock, but its electron temperature is not yet very high; predominantly, the effect of second-order bulk Comptonization in this zone leads to the formation of the tails in the emergent continuum; (3) the flow experiences the rather sharp braking in high-density layers of the shock wave up to a near-complete stop, photons gain the energy in this zone due to first-order and second-order bulk Comptonization, and by thermal Comptonization; the plasma can be heated to 10102020 keV and higher by means of the bulk heating mechanism there; (4) the matter falls, slowing down further, within a settling mound surrounded by the shock wave and the neutron star surface; as the matter settles, the electron temperature gradually approaches thermodynamically equilibrium one and becomes typically 1.51.544 times less than the temperature in the shock wave; (5) the plasma penetrates through the magnetic field beyond the channel near the base of the emitting region.
The assumption of LTE is a typical example of that which cannot actually work in the shock wave. Thus, Zhang et al. (2021) have used the Stefan-Boltzmann law to write the radiation temperature for the balance equation, which caused the obtaining of too low electron temperature in the shock, where the Compton heating and cooling must be modelled, with the bulk heating mechanism (Blandford & Payne, 1981b) must be taken into account. The velocity profiles in the solutions without allowing for modelling thermal Comptonization are not very sensitive to the assumption on the electron temperature, since the radiative diffusive and enthalpy fluxes must not be strongly dependent on it (Davidson, 1973; Arons et al., 1987). Corresponding transfer equation for the total radiation energy density is obtained from the transfer equation for the spectral radiation energy density under the assumption about the Compton equilibrium (Appendix A).
@@
The process of the going out of the plasma beyond the column needs to be modelled in detail in the framework of MHD simulations. Simple estimates were made by Lyubarskii & Syunyaev (1988). One can obtain that the mass growing above the neutron star crust can originate between outpourings a dense gas-dominated non-steady thermalized layer of the height $H\lesssim 2kT{R^{2}}/({GMm_{\rm p}})\approx 60~{\rm cm}$ (for the constant temperature taken for the upper part of the layer and set to be equal $6$ keV). This layer, or ‘the thermal mound’ (Becker, 1998), can be ignored in the calculations of the radiation transfer in the funnel, since its altitude is small and the energy exchange throughout the medium above the layer is significant.
The radiation energy density corresponding to the post-shock velocity $\sim 1/7$ of the pre-shock value cannot serve for the implementation of the bottom boundary condition, since this post-shock value of the velocity was obtained for an infinite shock wave in free space (e.g. Zel’dovich & Raizer 1967). The circumstances of X-ray pulsars (e.g. the boundedness of the settling mound by the neutron star surface, the transverse boundedness of the gas flow, the presence of a magnetic field) lead to $\gamma\sim 1$. The used value of $\gamma$ leads to $v(R)$ of $\sim 3\times 10^{8}$$4\times 10^{8}~{\rm cm~s^{-1}}$. A reasonable variation in $\gamma$ in both directions cannot change the results considerably (for $\gamma>0.99$ a steady solution does not often exist yet).
@@ 4.2 On the fit of observations
The high-energy tails, expanding mainly due to the influence of second-order bulk Comptonization, can become an obstacle to fit observed fallen quasi-power-law spectra with the developed model. The received flux in the tails is often relatively low (e.g. Kühnel et al. 2017). If future simulations taken into account angular photon redistribution and cyclotron resonances do not lead to better agreement with observations at usual problem parameters, this will mean that either the velocity at the top boundary may be distinguished noticeably from the free-fall velocity (and then the simplest solution is in the parametrization of the velocity at the top) or the realistic profiles of the shock are much more complicated than obtained here and must be found in the spectral radiative transfer simulations coupled with the MHD simulations. An experimental explanation may consist in the absence of the corresponding observational data. The obtained hardness of the spectrum taken near the energies $10$ keV and $1$ keV may exceed the observational value because of ignoring the reproduction of the radiation in the atmosphere of a neutron star, whose the albedo is characterised by a low-frequency cutoff (Postnov et al., 2015).
@@ 5 Conclusions
It has been found that the peak height of the the model accretion column, being calculated taking into account the specific form of the frequency- and polarization-dependent scattering cross sections, can depend nonlinearly on the accretion rate, in contrast to the nearly linear outputs obtained in the grey approximation. In the case of a sufficiently strong magnetic field, the saturation in the column peak height–accretion rate dependence takes place in the studied range of the accretion rate.
@@
The temperature structure should be calculated in general case with taking into account the influence of the induced Compton effect, which impact on the electron temperature within the settling mound under the condition of the sufficiently weak magnetic field and (in general case) on the existence of the steady-state solutions of the problem. The effects of second-order bulk Comptonization are important mainly for the formation of the spectra of emergent radiation, resulting in a smoothed power-law high-energy zone with a cutoff. I have discussed main directions for updating the results in the framework of more accurate models.
@@ Appendix A Equations for the total radiation energy density
Collecting (37) and (40) yelds (Wang & Frank, 1981)
@@
These equations are crudely valid while the flow is not too dense, and hence free-free processes are not important. Let us the second-order bulk Comptonization term is now taken into account. Without specification of the terms taken into account in the right-hand side of the momentum equation and of the way of the calculation of $T$, one has for the non-steady case (e.g. West et al. 2017b)

### v3

@@ 3.2 Results of the numerical calculations
The boundaries of the colour bars related to the 2D distributions approximately indicate the minimum and maximum values approached within the shown domains (for the velocity, these values fall within the corresponding boundary intervals). The difference between $\theta_{0}$, $\theta_{1}$ and $\theta_{2}$ in calculations and their values indicated above lies within the grid steps. Seediness at the front of the shock wave in the electron temperature distributions is the trace of the colour interpolation and is neither a manifestation of the scheme instabilities nor a property of the grid functions.
@@
It follows from the description above that, in simulations, the absolute values of PF increase with increasing accretion rate (Figs 13). That is, unsaturated Comptonization of the radiation of the ordinary mode can result in an emergent spectrum comparably close to the spectrum of the extraordinary mode (at least at $\dot{M}_{17}\lesssim 1$, Figs 1, 4, 5 and 6). It is especially clear in the case of the near-critical value of the accretion rate $\dot{M}_{17}=0.2$ ($r_{*}(r_{\rm up})=1.15$, $L_{\rm X,\,37}\approx 0.7$), see Fig. 6. The calculated PF in the range 112020 keV is in a satisfactory agreement with theoretical that of Burnard et al. (1991) and with characteristic observational values (e.g. Tsygankov et al. 2022; Long et al. 2023).
@@ 3.4 Effects of bulk Comptonization
In the absence of a magnetic field, first-order bulk Comptonization would be the reason of a shift in the maximum of the emergent spectrum towards higher energies relative to the characteristic energy $2.8kT_{0}$. This is a typical picture (see Psaltis & Lamb 1999, fig. 1). In a strong magnetic field this effect takes place as well. The emergent spectrum of the extraordinary mode is shifted to higher energies relative to the value of $kT_{0}$ mainly due to the influence of first-order bulk Comptonization, while the thermal effect of the shock wave is not very significant for the shift. The spectrum of the ordinary mode can be affected by the heating of the shock wave. If a high-energy hump in the emergent spectrum of the ordinary mode is formed, it usually corresponds to the energy $\sim 2kT_{\rm s}$, where $T_{\rm s}$ is the maximum electron temperature (approaching in the shock wave).
@@
The influence of the $m_{\rm e}v^{2}/3$ term can be illustrated by varying the velocity at the upper boundary. Consider a filled funnel. In Fig. 5 the structure, the emergent spectra and the polarization fraction are shown for $v(r_{\rm up})=1.3\times 10^{10}$ cm s-1 ($\dot{M}_{17}=1$, $L_{\rm X,\,37}\approx 1.7$). The decrease in $v(r_{\rm up})$ leads to the shift of the high-energy cutoff towards low energies, as expected. A considerable decrease in $v(r_{\rm up})$ (up to $\sim c/3$) demands a decrease in $\gamma$ for the existence of solutions in the framework of the model. Although such solutions contain a significantly cooler shock (up to $T_{\rm s}\approx 9$ keV at $T_{0}\approx 7$ keV for $\dot{M}_{17}=3$), the relatively softer spectral tail remains nevertheless rather hard, and the PF distribution is similar to that given in Fig. 2. The possibility of the distinction of $v(r_{\rm up})$ from the free-fall velocity was mentioned by Poutanen et al. (2013).
@@ 4.1 The model
The overall picture of matter breaking in the radiation-dominated regime is thus as follows. Five zones can be distinguished in the accretion channel above the magnetic pole: (1) the cold accretion flow falls near-freely in a radiation field of relatively low energy densities; (2) the flow begins to fluently brake in the shock, but its electron temperature is not yet very high; predominantly, the effect of second-order bulk Comptonization in this zone leads to the formation of the tails in the emergent continuum; (3) the flow experiences the rather sharp braking in high-density layers of the shock wave up to a near-complete stop, photons gain the energy in this zone due to first-order and second-order bulk Comptonization, and by thermal Comptonization; the plasma can be heated to 10102020 keV and higher by means of the bulk-heating mechanism there; (4) the matter falls, slowing down further, within a settling mound surrounded by the shock wave and the neutron star surface; as the matter settles, the electron temperature gradually approaches thermodynamically equilibrium one and becomes typically 1.51.544 times less than the temperature in the shock wave; (5) the plasma penetrates through the magnetic field beyond the channel near the base of the emitting region.
The assumption of LTE is a typical example of that which cannot actually work in the shock wave. Thus, Zhang et al. (2021) have used the Stefan-Boltzmann law to write the radiation temperature for the balance equation, which caused the obtaining of too low electron temperature in the shock, where the Compton heating and cooling must be modelled, with the bulk-heating mechanism (Blandford & Payne, 1981b) must be taken into account. The velocity profiles in the solutions without allowing for modelling thermal Comptonization are not very sensitive to the assumption on the electron temperature, since the radiative diffusive and enthalpy fluxes must not be strongly dependent on it (Davidson, 1973; Arons et al., 1987). Corresponding transfer equation for the total radiation energy density is obtained from the transfer equation for the spectral radiation energy density under the assumption about the Compton equilibrium (Appendix A).
@@
The process of the going out of the plasma beyond the column needs to be modelled in detail in the framework of MHD simulations. Simple estimates were made by Lyubarskii & Syunyaev (1988). One can obtain that the mass growing above the neutron star crust can originate between outpourings a dense gas-dominated non-steady thermalized layer of the height $Z\lesssim 2kT{R^{2}}/({GMm_{\rm p}})\approx 60~{\rm cm}$ (for the constant temperature taken to be equal $6$ keV). This layer, or ‘the thermal mound’ (Becker, 1998), can be ignored in the calculations of the radiation transfer in the funnel, since its altitude is small and the energy exchange throughout the medium above the layer is significant.
The radiation energy density corresponding to the post-shock velocity equal to $1/7$ of the pre-shock value cannot serve for the implementation of the bottom boundary condition, since this post-shock value of the velocity was obtained for a typical infinite shock wave (e.g. Zel’dovich & Raizer 1967). The circumstances of X-ray pulsars (e.g. the boundedness of the settling mound by the neutron star surface, the transverse boundedness of the gas flow, the presence of a magnetic field) lead to $\gamma\sim 1$. The used value of $\gamma$ leads to $v(R)$ of about $3\times 10^{8}$$4\times 10^{8}~{\rm cm~s^{-1}}$. A reasonable variation in $\gamma$ in both directions cannot change the results considerably (for $\gamma>0.99$ a steady solution does not often exist yet).
@@ 4.2 On the fit of observations
The high-energy tails, expanding mainly due to the influence of second-order bulk Comptonization, can become an obstacle to fit observed fallen quasi-power-law spectra with the developed model. The received flux in the tails is often relatively low (e.g. Kühnel et al. 2017). If future simulations taken into account angular photon redistribution and cyclotron resonances do not lead to better agreement with observations at usual problem parameters, this will mean that either the velocity at the top boundary may be distinguished noticeably from the free-fall velocity (and then the simplest solution is in the parametrization of the velocity at the top) or the realistic profiles of the shock are much more complicated than obtained here and must be found in the spectral radiative transfer simulations coupled with the MHD simulations. An experimental explanation may consist in the absence of the corresponding observational data. The obtained hardness of the spectrum taken near the energies $10$ keV and $1$ keV may exceed the observational value due to ignoring the reproduction of the radiation in the atmosphere of a neutron star, whose the albedo is characterised by a low-frequency cutoff (Postnov et al., 2015).
@@ 5 Conclusions
It has been found that the peak height of the model accretion column, being calculated taking into account the specific form of the frequency- and polarization-dependent scattering cross sections, can depend nonlinearly on the accretion rate, in contrast to the nearly linear outputs obtained in the grey approximation. In the case of a sufficiently strong magnetic field, the saturation in the column peak height–accretion rate dependence takes place in the studied range of the accretion rate.
@@
The temperature structure should be calculated in general case with taking into account the influence of the induced Compton effect, which impact on the electron temperature within the settling mound under the condition of the sufficiently weak magnetic field and (at the sufficient accretion rate) on the existence of the steady-state solutions of the problem. The effects of second-order bulk Comptonization are important mainly for the formation of the spectra of emergent radiation, resulting in a smoothed power-law high-energy zone with a cutoff. I have discussed main directions for updating the results in the framework of more accurate models.
@@ Appendix A Equations for the total radiation energy density
Collecting (37) and (40) yields (Wang & Frank, 1981)
@@
These equations are crudely valid while the flow is not too dense, and hence free-free processes are not important. Let the second-order bulk Comptonization term be now taken into account. Without specification of the terms taken into account in the right-hand side of the momentum equation and of the way of the calculation of $T$, one has for the non-steady case (e.g. West et al. 2017b)